Solving quadric equations by factorization

Find the real solutions of the following equations
1.(x^2+2x-14)^5=1
2.(x^2+7x-11)^8=1
1
Expert's answer
2020-12-24T18:12:32-0500

1.


"(x^2+2x-14)^5=1, x\in\R"

Let "x^2+2x-14=u, u\in\R"


"u^5-1=u^5-u^4+u^4-u^3+u^3-u^2+u^2-u+u-1"

"=(u-1)(u^4+u^3+u^2+u+1)"


"=(u-1)(u^4+u^3+\dfrac{1}{4}u^2+\dfrac{3}{4}u^2+u+1)"

"=(u-1)(u^2(u+\dfrac{1}{2})^2+\dfrac{1}{2}u^2+\dfrac{1}{4}u^2+u+1)"

"=(u-1)(u^2(u+\dfrac{1}{2})^2+\dfrac{1}{2}u^2+(\dfrac{1}{2}u+1)^2)"

"u^2(u+\dfrac{1}{2})^2+\dfrac{1}{2}u^2+(\dfrac{1}{2}u+1)^2>0, u\not=0, u\not=-\dfrac{1}{2}"

"u^2(u+\dfrac{1}{2})^2+\dfrac{1}{2}u^2+(\dfrac{1}{2}u+1)^2=1>0, u=0"

"u^2(u+\dfrac{1}{2})^2+\dfrac{1}{2}u^2+(\dfrac{1}{2}u+1)^2=\dfrac{1}{8}>0, u=-\dfrac{1}{2}"

Then


"u^2(u+\dfrac{1}{2})^2+\dfrac{1}{2}u^2+(\dfrac{1}{2}u+1)^2>0, u\in\R"

Hence


"u-1=0"


"x^2+2x-14-1=0"

"x^2+2x-15=0"

"(x+5)(x-3)=0"

"x+5=0\ or\ x-3=0"

"x_1=-5, x_2=3"

"\{-5, 3\}"


2.


"(x^2+7x-11)^8=1, x\in\R"

"((x^2+7x-11)^4-1)((x^2+7x-11)^4+1)=0"

Since "x\in \R," then "(x^2+7x-11)^4+1>0"



"(x^2+7x-11)^4-1=0"

"((x^2+7x-11)^2-1)((x^2+7x-11)^2+1)=0"

Since "x\in \R," then "(x^2+7x-11)^2+1>0"



"(x^2+7x-11)^2-1=0"

"((x^2+7x-11)-1)((x^2+7x-11)+1)=0"


"(x^2+7x-12)(x^2+7x-10)=0"

"x^2+7x-12=0\ or\ x^2+7x-10=0"

"x^2+7x-12=0"

"x^2+2(\dfrac{7}{2})x+(\dfrac{7}{2})^2-\dfrac{97}{4}=0"

"(x+\dfrac{7}{2})^2-\dfrac{97}{4}=0"

"x_1=-\dfrac{7}{2}-\dfrac{\sqrt{97}}{2}, x_2=-\dfrac{7}{2}+\dfrac{\sqrt{97}}{2}"

"x^2+7x-10=0"

"x^2+2(\dfrac{7}{2})x+(\dfrac{7}{2})^2-\dfrac{89}{4}=0"

"(x+\dfrac{7}{2})^2-\dfrac{89}{4}=0"

"x_3=-\dfrac{7}{2}-\dfrac{\sqrt{89}}{2}, x_4=-\dfrac{7}{2}+\dfrac{\sqrt{89}}{2}"

"\bigg\{-\dfrac{7+\sqrt{97}}{2},-\dfrac{7+\sqrt{89}}{2},\dfrac{-7+\sqrt{89}}{2},\dfrac{-7+\sqrt{97}}{2}\bigg\}"